OpenStudy (anonymous):

The height of an equilateral triangular prism increases by three units. The new volume is more than the original by how much? three more than the the area of the base three more than the length of the height three times the area of the base three times the length of the height

OpenStudy (anonymous):

help C:

hero (hero):

In order to solve this problem, you will need to know the following 1. Volume of an equilateral triangular prism 2. How to identify the dimensions of an equilateral triangular prism 3. How to apply the dimensions of an equilateral triangular prism to the formula for volume. 4. How to simplify algebraic expressions. Which of these concepts are you having trouble with?

OpenStudy (anonymous):

all of them . But i think it is A so ill just go with that i guesss

OpenStudy (anonymous):

??

hero (hero):

1. The volume of the triangular prism is V = Bh 2. The dimensions of the new Volume are B = 1/2b*h h = h + 3 3. After applying the dimensions to the formula for Volume of an equilateral triangular prism: \[V = \frac{bh(h+3)}{2} \] 4. Simplifying that we get: \[V= \frac{bh^2 + 3bh}{2}\]

OpenStudy (anonymous):

so it is C

OpenStudy (anonymous):

?

hero (hero):

It would probably be best if you used actual numbers. We know that the figure is an equilateral triangular prism, therefore, We can let the one side of the base = 4 and the height of the prism = 7 The height of the new prism will be 7 + 3 = 10

hero (hero):

So calculate the Volume of the old prism and the new prism, then figure out by how much the volume of the old prism has increased.

OpenStudy (anonymous):

three times area of the base

OpenStudy (anonymous):

good Question. :)

hero (hero):

Did you calculate that, or are you just guessing? I don't want to have to do all the work here. I'd rather see you calculate the Volumes based on what I have posted above and let me know what you got.

OpenStudy (anonymous):

ill just guess , sorry for wasting your time ? :c